OpenStudy (anonymous):

Give the 5th and 10th term for the sequence formula below. tn = 1 + 6n Answer 30, 60 31, 61 15, 16 11, 16

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

t5=1+6*5=31 t10=1+6*10=61

OpenStudy (anonymous):

got it?